Find the length and width of rectangle. Let x+7 be length and x be the width. 198 in^2 is the area.
Please help me solve this problem!

Answer by aswathytony(47)


given the length as (x + 7)
width x
then Area = l*w
= (x+7) * x ..........(1)
given area = 198 squ .units

i.e . (x + 7) x = 198
x^2 + 7x = 198
x^2 +7x -198 =0.
x^2 + 18 x -11x -198 =0
x ( x +18 ) -11 (x + 18 ) = 0
(x - 11) ( x +18 ) = 0
x = 11 or x = -18
since x is width possible value is x = 11 .
l = x + 7 = 11 +7 = 18
i.e. length = 18 units and width = 11 units